Grade 12Magical Mathematics[Interesting Approach]

How many pairs of natural numbers are there, the difference of whose squares is 45?
a) 3 c) 4
b) 5 d) 6

Divnoor singh
10 Years agoGrade 12

1 Answer

Harsh Patodia
ApprovedApproved Tutor Answer10 Years ago
a​2 - b​2 =(a-b) (a+b) = 45 = 1*45
= 3*15
= 9*5
just solve these cases and seefrom which natural number are coming.
for eg a-b=5 a+b=9
from this a=7and b=2